OpenStudy (anonymous):

What is the solution of (square root 5)^3 Option A: 5 3/2 Option B: 5 2/3 Option C: 5 1/3 Option D: 5 1/6

OpenStudy (anonymous):

\[\huge \sqrt[2]{5^3}=5^{\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

or more commonly written as \[5\sqrt{5}\]

OpenStudy (anonymous):

Thanks!